# Space Group Informed Transformer for Crystalline Materials Generation

### v1

@@ Abstract
We introduce CrystalFormer, a transformer-based autoregressive model specifically designed for space group-controlled generation of crystalline materials. The space group symmetry significantly simplifies the crystal space, which is crucial for data and compute efficient generative modeling of crystalline materials. Leveraging the prominent discrete and sequential nature of the Wyckoff positions, CrystalFormer learns to generate crystals by directly predicting the species and locations of symmetry-inequivalent atoms in the unit cell. Our results demonstrate that CrystalFormer matches state-of-the-art performance on standard benchmarks for both validity, novelty, and stability of the generated crystalline materials. Our analysis also shows that CrystalFormer ingests sensible solid-state chemistry information from data for generative modeling. The CrystalFormer unifies symmetry-based structure search and generative pre-training in the realm of crystalline materials. The simplicity, generality, and flexibility of CrystalFormer position it as a promising architecture to be the foundational model of the entire crystalline materials space, heralding a new era in materials modeling and discovery.
@@ I Introduction
Machine learning methods are playing an increasingly important role in material discovery, complementing conventional computational approaches [1, 2]. Generative machine learning, in particular, has been a promising step for matter inverse design [3, 4] which goes beyond machine learning accelerated structure search [5] and property screening [6]. Generative models learn the underlying distribution of training data and generated new samples from the learned distribution. In addition, the generation process can also be controlled by prompts such as desired material properties or experiment observations. Amazing programming abilities of generative models have been demonstrated in large language model [7], text-to-image generation [8, 9], and protein design [10].
It is anticipated that generative model-based approaches will introduce groundbreaking changes to the traditional workflows of material discovery. A generative pretrained foundation model for crystalline materials is a key step towards such lofty goal. However, despite of intensive efforts [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], the current generative models for crystalline materials fall short to match the success of other domains. Simply scaling the compute and model size of the current crystal generative model may not be enough because the amount of high-quality data for crystalline materials is much less than compared to language and image domains. Therefore, leveraging the inherent inductive biases specific to crystalline structures for more data-efficient generative modeling is essential, as has been pursued in some of recent works [12, 23, 24, 25, 26].
@@
The space group symmetry is arguably the most important inductive bias in the modeling of crystalline materials, which is the joint outcome of the rotational and translational symmetry in space. There are in total 230 space groups [29] for three-dimensional crystal structures. Nature exhibits a preference for symmetric crystal structures, a tendency that may be attributed to the symmetry inherent in the interatomic interactions, which, in turn, are governed by the fundamental forces acting between elementary particles. As a result, the appearance of crystalline materials in the the first and the least symmetric space group is rare [30], with many instances potentially being misclassified. [31].
@@
Space group symmetry imposes significant constraints on a crystal. At its core, the space group identifies the crystal system to which a crystal belongs, thereby limiting the permissible values for the lattice parameters that define the dimensions and angles of the crystal’s unit cell. Moreover, the symmetry operations associated with a given space group ensure that identical atoms are consistently mapped onto each other across the crystal, preserving its structural integrity. This requirement enforces strict conditions regarding the types of chemical elements present, their specific locations within the crystal, and the total number of atoms in each unit cell. A key concept in understanding these constraints is the Wyckoff positions, which delineate unique areas within a unit cell that are defined by the symmetry operations of the crystal’s space group. These positions are represented as fractional coordinates, enabling precise definition relative to the unit cell’s axes. For example, Fig. 1(a) shows the Wyckoff positions for the space group (No. 167). The Wyckoff positions are labeled by letters in the alphabet, starting from special points in the bottom to general positions in the top. The multiplicity counts the number of equivalent positions connected by the space group symmetry operations. All of them should be occupied by the same type of atoms to uphold the space group symmetry. For example, the top row of the table in Fig. 1(a) shows a general position that can be mapped to 36 positions under the symmetry operations of the space group.
@@
Nature tends to place atoms in those special Wyckoff positions at the bottom of the table. For example, we highlight the occupied Wyckoff positions of calcite (CaCO) crystal in Fig. 1, associated with the space group. One sees that the Wyckoff letter ’6a’ and ’6b’ deterministicly define the locations of the carbon and calcium atoms within the unit cell. In addition, it follows that , and as the space group belongs to the trigonal crystal system. Ultimately, despite having 30 atoms in the unit cell, there are only three continuous degrees of freedom for the CaCO structure: the coordinate of oxygen atom and the lattice constants and . All other information about the crystal structure can be specified via discrete data such as the Wyckoff letters and chemical species.
@@
The prominent discrete and sequential features illustrated in Figure 1 are ubiquitous in crystalline materials. The Wyckoff positions not only specify possible locations of atoms in the unit cell, their associated multiplicities also put strong constraints on the number of atoms. Therefore, space group symmetry significantly reduces the degrees of freedom of crystalline materials. Failing to exploit this information in generative modeling not only renders learning inefficient, it also severely impairs the generalization ability of the model. For example, the performance of the generative model quickly deteriorates as the number of atoms increases due to it is challenging to generate highly symmetric crystal structures [16].
@@
In this paper, we introduce CrystalFormer, an autoregressive transformer for generative modeling of crystalline materials. CrystalFormer models the joint probability distribution of Wyckoff positions, chemical species, and lattice parameters of crystals with a given space group. By treating the Wyckoff positions as the first class citizen in the model, CrystalFormer seamlessly integrates the space group symmetry into crystal probabilistic modeling. The space group-informed transformer exploits the fundamental inductive bias of crystals to greatly constrain and simplify the generation process. As shown in Table 1, explicit modeling of the Wyckoff positions greatly reduces the space of crystalline materials. We benchmark the CrystalFormer model on standard crystal generation tasks and show it exhibits great efficiency and generalization ability compared to existing models. Finally, we discuss its potential applications in novel materials discovery. We have released the codes and trained model at [32].
@@ II CrystalFormer
To exploit the space group symmetry of the crystal, we focus on the Wyckoff positions of symmetry-inequivalent atoms. Wyckoff letters follows the alphabetical order, where ”a” stands for the positions with the highest order of symmetry for the given space group. Later letters in the alphabet indicate more general positions with reduced site symmetries. Note that the information of the space group number and Wyckoff letter fully determine the multiplicities. In cases where the atom positions are fully fixed by the Wyckoff letter, we will also consider the remaining fractional coordinates, e.g. the -coordinate of the oxygen atoms in the CaCO example shown in Fig. 1, To generate crystals, one samples the Wyckoff letter, chemical element, and fractional coordinates of each atom sequentially. The sampling procedure starts from special higher symmetry sites with smaller multiplicities and then goes on to general lower symmetry regions with larger multiplicities.
With these considerations, we define a crystal data as . Here are Wyckoff letters and are chemical species. Here, stands for the number of symmetrically inequivalent atoms in the conventional unit cell. For example, as shown in Fig. 1(b) one has for CaCO. Explicitly including the Wyckoff letter in the generative modeling is the key of the present work. Next, are the fractional coordinates of symmetrically inequivalent atoms. Lastly, denotes the lattice parameters of the conventional unit cell of the material.
The central quantity to focus on is the conditional probability of a crystal given the space group number : . Since the space group is a fundamental characterization for crystalline materials, is a key control variable that greatly simplifies the distribution over the entire crystal materials space. In practical applications, the space group can be considered separately or even predicted [33, 34].
@@
At first sight, it may appear unnatural to employ an autoregressive model for crystals since there is no obvious order for atoms in the unit cell. However, the sequential nature of Wyckoff positions suggests a natural way to arrange symmetrically inequivalent atoms in an alphabetical order of the Wyckoff letters. During sampling, the autoregressive model places the atoms one by one into the unit cell, starting from the special position with high symmetry to the general position with the lowest symmetry, see Fig. 1(c).
@@
We model the conditional probability of the Wyckoff letters and chemical species as categorical distributions. A mixture of von-Mises distribution is employed to model the conditional probability of the factional coordinates . For Wyckoff positions with multiplicities greater than one, we only consider the first of fractional coordinates that appear in the international tables for crystallography [35]. Lastly, we model the conditioned distribution of lattice parameters as a Gaussian mixture model.
@@
We build CrystalFormer using the vanilla decoder-only transformer [36] to model the probability distribution of crystalline materials Eq. (1). The transformer-based autoregressive model is a favorable choice to capture these prominent discrete characteristics of crystals. We feed atomistic features into an autoregressive transformer. The space group number is the first input to the transformer. The remaining inputs alternate between the Wyckoff letter, chemical species, and fractional coordinates of each atom. One can go through the table of Fig. 1(b) in a raster order to collect these atomistic features. We feed vector embeddings of the space group number, Wyckoff letter, and the chemical species input to the CrystalFormer. In particular, we concatenate the vector embedding of to all other inputs since it is the key control variable for the crystal generation. In addition, we have also provided the multiplicity of each Wyckoff position as an additional feature. The multiplicity can be easily inferred from the space group and the Wyckoff letters. We feed the fractional coordinates as Fourier features into the transformer so that the model preserves the periodicity of the unit cell [37, 13] We pad the atom sequence up to a maximum length and take the shifted output as parameters of the conditional probability distribution Eq. (1), see Fig. 1(b). At the location of the first padding atom, we predict the lattice parameters.
@@
We implement a number of constraints in the model to further reduce the phase space of the model. First, the Wyckoff letters should be valid for the given space group. For example, for the space group number 167 the Wyckoff letters go from ’a’ up to ’f’. Second, we require that the Wyckoff letters follow alphabetical order in the sequence 111We follow the convention that capital letters appear after lower case letters. This handles the edge case of the No. 47 space group whose Wyckoff positions reach ’A’. This rule also demands that the padded atoms appear at the end of the sequence with Wyckoff letter ’X’.. Lastly, the Wyckoff positions with no continuous degree of freedom (such as ’a’, ’b’, and ’d’ positions in the space group 167) can only be occupied once. Those constraints are implemented by tuning the logit biases of Wyckoff letters [39, 40].
@@
The design of CrystalFormer focuses mostly on the space group symmetries since we believe they are the most crucial ones for crystalline materials. This design decision significantly impacts the treatment of other symmetries. First, it is often possible to place the origin of the unit cell at the inversion center of the specified space group. The chosen origin naturally fixes the continuous translation invariance of fractional coordinates. Second, by only considering symmetry-inequivalent atoms and labeling them with Wyckoff letters, one fixes most of the permutation invariance in the representation. For those Wyckoff positions with continuous degrees of freedom, there may be multiple symmetry-inequivalent atoms with the same Wyckoff letters. We arrange these atoms according to the lexicographic order of fractional coordinates [41] in the sequence. Note that in a crystal environment, the same type of atoms occupying different Wyckoff positions could be regarded as distinguished particles as they generally have different site symmetry. Lastly, the periodicity of the fractional coordinates is respected since they are treated as variables under the von-Mises distribution.
@@
To sample crystals from the CrystalFormer, one needs to specify a space group number and a number of possible chemical elements. The CrystalFormer samples the atoms one by one, starting from more symmetric specific positions with lower multiplicities till less symmetric general positions with larger multiplicities. We use the information of the space group and Wyckoff letter to control the sampling of fractional coordinates. By applying the symmetry projection to the sampled fractional coordinate, one ensures the generated fractional coordinates are compatible with the Wyckoff positions. One can also mask out the logits of chemical species so that only a number of selected elements will be sampled. The number of symmetrically inequalivent atoms may fluctuate in the sampling procedure. Once one has sampled a padding atom, the model predicts the lattice parameters under the space group constraint. Moreover, we introduce a temperature parameter in the sample distribution . With we will draw samples from a sharper distribution, while gives more diversity in the generated samples.
III Results and Evaluations

### v2

@@ Abstract
We introduce CrystalFormer, a transformer-based autoregressive model specifically designed for space group-controlled generation of crystalline materials. The incorporation of space group symmetry significantly simplifies the crystal space, which is crucial for data and compute efficient generative modeling of crystalline materials. Leveraging the prominent discrete and sequential nature of the Wyckoff positions, CrystalFormer learns to generate crystals by directly predicting the species and locations of symmetry-inequivalent atoms in the unit cell. We demonstrate the advantages of CrystalFormer in standard tasks such as symmetric structure initialization and element substitution compared to conventional methods implemented in popular crystal structure prediction software. Moreover, we showcase the application of CrystalFormer of property-guided materials design in a plug-and-play manner. Our analysis shows that CrystalFormer ingests sensible solid-state chemistry knowledge and heuristics by compressing the material dataset, thus enabling systematic exploration of crystalline materials. The simplicity, generality, and flexibility of CrystalFormer position it as a promising architecture to be the foundational model of the entire crystalline materials space, heralding a new era in materials modeling and discovery.
Contents
@@
- II CrystalFormer
- III CrystalFormer learns chemical intuition by compressing materials database
- IV Applications
- V Related works
- VI Outlook
- A More details of CrystalFormer
- B Validity and novelty of generated samples
- C Discovered crystal samples with symmetric structure initialization
- D Discovered crystal samples with element substitution
- E Details of plug-and-play materials design
I Introduction
Machine learning methods are playing an increasingly important role in material discovery, complementing conventional computational approaches [1, 2]. Generative machine learning, in particular, has been a promising step for matter inverse design [3, 4] which goes beyond machine learning accelerated structure search [5] and property screening [6]. Generative models learn the underlying distribution of training data and generate new samples from the learned distribution. In addition, the generation process can also be controlled by conditions such as desired material properties or experiment observations. Amazing programming abilities of generative models have been demonstrated in large language model [7], text-to-image generation [8, 9], and protein design [10].
It is anticipated that generative model-based approaches will introduce groundbreaking changes to the traditional workflows of material discovery. A generative pre-trained foundation model for crystalline materials is a key step towards such a lofty goal. However, despite intensive efforts [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], the current generative models for crystalline materials fall short to match the success of other domains. Simply scaling the compute and model size of the current crystal generative model may not be feasible because the amount of high-quality data for crystalline materials is much less than compared to language and image domains. Therefore, leveraging the inherent inductive biases specific to crystalline structures for more data-efficient generative modeling is essential, as has been pursued in some of recent works [23, 24, 25, 26].
@@ I Introduction
The space group symmetry due to the joint outcome of the rotational and translational symmetry in space is arguably the most important inductive bias in the modeling of crystalline materials. There are in total 230 space groups [29] for three-dimensional crystal structures. Nature exhibits a preference for symmetric crystal structures, a tendency that may be attributed to the symmetry inherent in the interatomic interactions, which, in turn, are governed by the fundamental forces acting between elementary particles. As a result, the appearance of crystalline materials in the first and the least symmetric space group is rare [30], with many instances potentially even being misclassified [31]. Failing to match the space group distribution of nature in machine learning-generated materials is regarded as a matter of serious concern [32].
@@
Space group symmetry imposes significant constraints on a crystal. First of all, the space group identifies the crystal system to which a crystal belongs, thereby limiting the permissible values for the lattice parameters that define the length and angles of the crystal’s unit cell. Moreover, the symmetry operations associated with a given space group ensure that symmetry equivalent atoms are consistently mapped among themselves in the crystal. This requirement enforces strict conditions regarding the types of chemical elements present, their specific locations within the crystal, and the number of each chemical species in the unit cell. A key concept to express these constraints is the Wyckoff positions, which delineate unique areas within a unit cell that are defined by the symmetry operations of the crystal’s space group. These positions are represented as fractional coordinates, enabling precise definition relative to the unit cell’s axes. For example, Fig. 1(a) shows the Wyckoff positions for the space group (No. 167). The Wyckoff positions are labeled by letters in the alphabet, starting from special points in the bottom to general positions in the top. The multiplicity counts the number of equivalent positions connected by the space group symmetry operations. All of them should be occupied by the same type of atoms to uphold the space group symmetry. For example, the top row of the table in Fig. 1(a) contains the general position that can be mapped to 36 positions under the symmetry operations of the space group.
@@
Nature tends to place atoms in those special Wyckoff positions at the bottom of the table. For example, we highlight the occupied Wyckoff positions of calcite (CaCO3) crystal in Fig. 1, associated with the space group. One sees that the Wyckoff letter ’6a’ and ’6b’ deterministically define the locations of the carbon and calcium atoms within the unit cell. In addition, it follows that , and as the space group belongs to the trigonal crystal system. Ultimately, despite having 30 atoms in the unit cell, there are only three continuous degrees of freedom for the CaCO3 structure: the x-coordinate of oxygen atom and the lattice constants and . All other information about the crystal structure can be specified via discrete data such as the Wyckoff letters and chemical species.
@@
The prominent discrete and sequential features illustrated in Figure 1 are ubiquitous in crystalline materials. The Wyckoff positions not only specify possible locations of atoms in the unit cell, but their associated multiplicities also put strong constraints on the number of atoms. Therefore, space group symmetry significantly reduces the degrees of freedom of crystalline materials. Failing to exploit this information in generative modeling not only renders learning inefficient, it also severely impairs the generalization ability of the model. For example, the performance of the generative model quickly deteriorates as the number of atoms increases due to it is challenging to generate highly symmetric crystal structures [16]. On the other hand, statistical analysis shows that the Wyckoff sequences of known inorganic compounds [33] are far from being exhausted, implying there are statistical correlations to be exploited to compress the materials database.
@@
In this paper, we introduce CrystalFormer, an autoregressive transformer for generative modeling of crystalline materials. CrystalFormer models the joint probability distribution of Wyckoff positions, chemical species, and lattice parameters of crystals with a given space group. By treating the Wyckoff positions as the first class citizen in the model, CrystalFormer seamlessly and rigorously integrates the space group symmetry into crystal probabilistic modeling. As shown in Table 1, explicit modeling of the Wyckoff positions greatly reduces the space of crystalline materials. The space group-informed transformer exploits this fundamental inductive bias to greatly simplify the learning and generation of crystals.
The organization of the present paper is as follows. Section II introduces the model architecture, training, and sampling of the CrystalFormer. Section III reveals the chemical intuition encoded in the trained model by inspecting its weights and generated crystal samples. Such inspection also builds up an understanding of the strength of the model. Section IV demonstrates applications CrystalFormer for symmetric structure initialization and element substitution, where it exhibits great efficiency and generality compared to existing approaches. Furthermore, we showcase CrystalFormer’s ability for property-guided materials design in a plug-and-play manner. Section V puts the key contribution of present work in the broad context of crystal generative models. Finally, Section VI outlooks for possible extension and future of the present line of research. We have released the codes and trained model at [34].
@@ II CrystalFormer
We will introduce the design, training, and sampling of the CrystalFormer model.
II.1 Model
To exploit the space group symmetry of the crystal, we focus on the Wyckoff positions of symmetry-inequivalent atoms. Wyckoff letters follow the alphabetical order, where ”a” stands for the positions with the highest order of symmetry for the given space group. Later letters in the alphabet indicate more general positions with reduced site symmetries. Note that the information of the space group number and Wyckoff letter fully determine the multiplicities. In cases where the atom positions are not fully fixed by the Wyckoff letter, we will also consider the remaining fractional coordinates, e.g. the -coordinate of the oxygen atoms in the CaCO3 example shown in Fig. 1. To generate crystals, one samples the Wyckoff letter, chemical element, and fractional coordinates of each atom sequentially. The sampling procedure starts from special higher symmetry sites with smaller multiplicities and then goes on to general lower symmetry regions with larger multiplicities.
With these considerations, we define a crystal data as . Here are Wyckoff letters and are chemical species. Here, stands for the number of symmetrically inequivalent atoms in the conventional unit cell. For example, as shown in Fig. 1(b) one has for CaCO3. Explicitly including the Wyckoff letter in the generative modeling is the key of the present work. Next, are the fractional coordinates of symmetrically inequivalent atoms. Lastly, denotes the lattice parameters of the conventional unit cell of the crystal.
The central quantity to focus on is the conditional probability of a crystal given the space group number : . Since the space group is a fundamental characterization for crystalline materials, is a key control variable that greatly simplifies the distribution over the entire crystal materials space. In practical applications of crystal structure prediction and material design, the space group can either be considered separately as a control variable or predicted based on material composition [35, 36, 37, 38].
@@ II.1 Model
At first sight, it may appear unnatural to employ an autoregressive model for crystals since there seems to be no obvious order for atoms in the unit cell. However, the sequential nature of Wyckoff positions suggests a natural way to arrange symmetrically inequivalent atoms in an alphabetical order of the Wyckoff letters. Following this key observation, we represent crystal data as sequences of space groups, Wyckoff letters, chemical species, and fractional coordinates of each symmetrically inequivalent atom. Together with the information lattice parameters, such sequence fully characterizes the compositional and structural information of crystalline material. Since statistical analysis reveals that anions are in less symmetric positions than cations for inorganic crystals [30], one would expect that anion atoms will typically appear after cation atoms in such a sequence. For example, CaCO3 is represented as a string ”167-a-C-0-0-1/4-b-Ca-0-0-0-e-O-0.257-0-1/4”. Autoregressive sampling of such a string means the model generates the crystal by placing the atoms sequentially into the unit cell, starting from the special position with high site symmetry to the general position with the lowest site symmetry, see Fig. 1(d).
@@
We model the conditional probability of the Wyckoff letters and chemical species as categorical distributions. On the other hand, we model the conditional probability of the factional coordinates as a mixture of von Mises distribution for continuous periodic variables. For Wyckoff positions with multiplicities greater than one, we only consider the first of fractional coordinates that appear in the international tables for crystallography [39]. Lastly, we model the conditioned distribution of lattice parameters as a Gaussian mixture model.
@@
We build CrystalFormer, an autoregressive transformer [40] to model the space group conditioned-probability distribution of crystalline materials Eq. (1). The space group number is the first input to CrystalFormer. The remaining inputs are the Wyckoff letter, chemical species, and fractional coordinates of each atom. One can go through the table of Fig. 1(b) in a raster order to collect these atomistic features. We feed vector embeddings of the space group number, Wyckoff letter, and the chemical species input to the CrystalFormer. In particular, we also concatenate the vector embedding of to all other inputs since it is the key control variable for the crystal generation. Moreover, we have also provided the multiplicity of each Wyckoff position as an additional feature. The multiplicity can be easily inferred from the space group and the Wyckoff letters. We feed the fractional coordinates as Fourier features into the transformer so that the model preserves the periodicity of the unit cell [41, 13]. We pad the atom sequence up to a maximum length and treat the output as parameters of the conditional probability distribution Eq. (1), see Fig. 1(b). At the location of the first padding atom, we predict the lattice parameters.
@@
We implement a number of constraints in the model to further reduce its phase space. First, the Wyckoff letters should be valid for the given space group. For example, for the space group (No. 167) the Wyckoff letters go from ’a’ up to ’f’. Second, we require that the Wyckoff letters follow alphabetical order in the sequence 111We follow the convention that capital letters appear after lower case letters. This handles the edge case of the space group (No. 47) whose Wyckoff letters used up 26 lowercase letters and reached ’A’ for the generic position. In addition, we use the letter ’X’ to indicate the Wyckoff position of padding atoms that appear at the end of the sequence, see e.g. of Fig. 1(b).. Lastly, the Wyckoff positions with no free fractional coordinates (such as ’a’, ’b’, and ’d’ positions in the space group) can only be occupied once. Those constraints are implemented by setting the logit biases of Wyckoff letters to mask out invalid sequences [43, 44].
@@
The design of CrystalFormer focuses mostly on the space group symmetries which we believe to be the most important inductive bias for crystalline materials. This design decision significantly impacts the treatment of other symmetries. First, it is often possible to place the origin of the unit cell at the inversion center of the specified space group. The chosen origin naturally fixes the continuous translation invariance of fractional coordinates. Second, by only considering symmetry-inequivalent atoms and labeling them with Wyckoff letters, one fixes most of the permutation invariance in the representation. For those Wyckoff positions with continuous degrees of freedom, there may be multiple symmetry-inequivalent atoms with the same Wyckoff letters. We arrange these atoms according to the lexicographic order of fractional coordinates [45] in the sequence. Note that in a crystal environment, the same type of atoms occupying different Wyckoff positions could be regarded as distinguished particles as they generally have different site symmetry. Lastly, the periodicity of the fractional coordinates is respected in CrystalFormer since they are treated as periodic variables under the von Mises distribution.
II.2 Training